OpenStudy (loser66):

Let m, n be positive integers, prove that m!n! is a divisor or (m+n)! Please, help

OpenStudy (mom.):

*of

Parth (parthkohli):

\[\frac{(m+n)!}{m!n!}\]This expression happens to be the number of ways you can divide \(m+n\) objects into two groups of \(m\) and \(n\) objects. Which is obviously integral.

Parth (parthkohli):

Or you could think of it as the coefficient of \(x^m y^n\) in the expansion of \((x+y)^{m+n}\)

OpenStudy (loser66):

This is my attempt but I doubt myself when it is so easy. WLOG m < n < m+n hence m ! is a divisor of n! and n! is a divisor of (m+n)!

Parth (parthkohli):

Eh, you want to take that route. It's pretty clear that your expression is \(\large \binom{m+n}{m}\) but still...\[\frac{(m+n)!}{m!n!}\]\[= \frac{(m+1)(m+2)(m+3) \cdots (m+n)}{1\cdot 2 \cdots \cdot n}\]Pigeonhole Principle tells us that in a series of consecutive \(n\) numbers, we have at least one that is divisible by \(n\), at least one divisible by \(n-1\), and so on.

OpenStudy (loser66):

Hold on, that is what we have to prove.

OpenStudy (loser66):

Only thing I think I can expand is (m+n)! = (m+n) (m+n -1) (m+n-2) .........(m+n -m) (n-1) (n-2).....2*1

OpenStudy (loser66):

the last part is n!

ganeshie8 (ganeshie8):

combinatoric method that PK has shown earlier is a legitimate proof

OpenStudy (loser66):

One more thing, my prof gives me hint: You may use the fact that \( [x] +[y]\leq [x+y]\) where [ ] is floor function

ganeshie8 (ganeshie8):

If you have \(m+n\) people in your class, then the number of ways of choosing \(m\) people for basketball team is given by \(\dbinom{m+n}{m}\). Clearly, the number of choices cannot be a fraction.

ganeshie8 (ganeshie8):

I know that proof using floor funciton, one sec...

ganeshie8 (ganeshie8):

consider the prime factorization of \(n!\). do you remember how to get the exponent of a prime \(p\) ?

ganeshie8 (ganeshie8):

for example, what is the exponent of \(3\) in the prime factorization of \(20!\) ?

OpenStudy (loser66):

yes, it is \[\lfloor x/p\rfloor+\lfloor x/p^2\rfloor+.....\]

ganeshie8 (ganeshie8):

good, remember that, we're going to need that in the proof using floor functions

OpenStudy (loser66):

Show me, please

ganeshie8 (ganeshie8):

maybe below could be an even better start : for each prime factor \(p\) of the denominator term, \(m!n!\), we have : \[\left[ \dfrac{\color{red}{m}+\color{blue}{n}}{p^k}\right]\ge \left[\dfrac{\color{red}{m}}{p^k}\right]+\left[\dfrac{\color{green}{n}}{p^k}\right]\]

ganeshie8 (ganeshie8):

fine with that ? i have just used the given hint, nothing fancy..

OpenStudy (loser66):

Go ahead, I just don't see the link yet but pretty sure it will pop out next.

OpenStudy (loser66):

I got the equation, next?

OpenStudy (loser66):

the exponent of prime p on (m+n)!

ganeshie8 (ganeshie8):

For each prime factor \(p\) of the denominator term, \(m!n!\), we have : \[\left[ \dfrac{\color{red}{m}+\color{blue}{n}}{p^k}\right]\ge \left[\dfrac{\color{red}{m}}{p^k}\right]+\left[\dfrac{\color{green}{n}}{p^k}\right]\] It follows : \[\sum\limits_{k\ge 1}\left[ \dfrac{\color{red}{m}+\color{blue}{n}}{p^k}\right]\ge \sum\limits_{k\ge 1}\left[\dfrac{\color{red}{m}}{p^k}\right]+\sum\limits_{k\ge 1}\left[\dfrac{\color{green}{n}}{p^k}\right]\]

ganeshie8 (ganeshie8):

what does left hand side expression, \(\sum\limits_{k\ge 1} \left[ \dfrac{\color{red}{m}+\color{blue}{n}}{p^k}\right]\), represent ?

OpenStudy (loser66):

The total exponent of p on (m+n)! where p is a prime factor of (m+n)!

ganeshie8 (ganeshie8):

good, what about the right hand expression, \(\sum\limits_{k\ge 1}\left[\dfrac{\color{red}{m}}{p^k}\right]+\sum\limits_{k\ge 1}\left[\dfrac{\color{green}{n}}{p^k}\right]\) what does that represent ?

OpenStudy (loser66):

oh, the same, for the first one it is exponent of p on m! , second is on n!

ganeshie8 (ganeshie8):

right, when you add both exponents, you get the exponent of the denominator, \(m!n!\), yes ?

ganeshie8 (ganeshie8):

so, the right hand side represents the exponent of \(p\) in the prime factorization of \(m!n!\)

OpenStudy (loser66):

And since they are exponent , hence the right hand side becomes p^m p^n

OpenStudy (loser66):

Got you. Thank you so much.

ganeshie8 (ganeshie8):

what do you mean by : `hence the right hand side becomes p^m p^n `

OpenStudy (loser66):

oh, just too exciting when understanding it. hehehe... I know it is not the correct notation,

OpenStudy (loser66):

the first sum is \(p^{r}\) which is a prime factor of m! the second sum is \(p^{s}\) which is a prime factor of n! hence \(p^{r+s}\) is a prime factor of m! n! right?

OpenStudy (loser66):

And this prime is < the same prime factor of (m+n)! from the left hand side That shows m!n! is a factor of (m+n)! , right?

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (loser66):

Thanks again.

ganeshie8 (ganeshie8):

np